A copper alloy that is 40% copper is to be combined with a copper alloy that is 80% copper to produce 120 kilograms of an alloy that is 70% copper. How many kilograms of each alloy must be used?
Answer by stanbon(75887)

A copper alloy that is 40% copper is to be combined with a copper alloy that is 80% copper to produce 120 kilograms of an alloy that is 70% copper. How many kilograms of each alloy must be used?
-------------------
Equation:
copper + copper = copper
0.40x + 0.80(120-x) = 0.70*120
---
Multiply thru by 100 to get:
40x + 80*120 - 80x = 70*120
------
-40x = -10*120
x = (1/4)120
x = 30 kg (amt. of 40% alloy needed in the mixture.
=================
Cheers,
Stan H.